OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2.

OpenStudy (anonymous):

\[\frac{6+3}{x-10}=-\frac{3}{2}\] solve for \(x\)

OpenStudy (anonymous):

\[\frac{9}{x-10}=-\frac{3}{2}\] \[-2\times 9=3(x-10)\] etc

OpenStudy (anonymous):

\[-18=-30x\]